# Multi-Margin Loss: Proposal and Application in Recommender Systems

### v1

@@ Abstract.
Recommender systems guide users through vast amounts of information by suggesting items based on their predicted preferences. Collaborative filtering-based deep learning techniques have regained popularity due to their straightforward nature, relying only on user-item interactions. Typically, these systems consist of three main components: an interaction module, a loss function, and a negative sampling strategy. Initially, researchers focused on enhancing performance by developing complex interaction modules. However, there has been a recent shift toward refining loss functions and negative sampling strategies. This shift has led to an increased interest in contrastive learning, which pulls similar pairs closer while pushing dissimilar ones apart. Contrastive learning involves key practices such as heavy data augmentation, large batch sizes, and hard-negative sampling, but these also bring challenges like high memory demands and under-utilization of some negative samples. The proposed Multi-Margin Loss (MML) addresses these challenges by introducing multiple margins and varying weights for negative samples. This allows MML to efficiently utilize not only the hardest negatives but also other non-trivial negatives, offering a simpler yet effective loss function that outperforms more complex methods, especially when resources are limited. Experiments on two well-known datasets demonstrated that MML achieved up to a 20% performance improvement compared to a baseline contrastive loss function when fewer number of negative samples are used.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Neural networks
@@ 1. Introduction
Figure 1. Three major components for a deep-learning based CF model: (i) interaction module/encoder, (ii) loss function and (iii) negative sampling strategy.
Given a representation space, user interacts with an item (e.g., purchases a product), shown with $p$, and from the remaining negative items $n$, four negatives, $n_{1}-n_{4}$, are sampled. Then, the sampled negative items, positive item and user representations are used by the interaction encoding and loss function modules for the computations.
@@
Numerous web applications, such as YouTube, Netflix, Amazon, Spotify employ recommendation systems to assist users in navigating the overwhelming amount of content (Park et al., 2023). These systems suggest items, e.g., videos, products, and music tracks, by estimating user preferences (Massa and Avesani, 2007; Tavakolifard and Almeroth, 2012). In the literature, recommender systems employ various approaches, from traditional methods like collaborative filtering, content-based filtering, matrix factorization(Pan et al., 2008; Ye et al., 2010; Li et al., 2015; He et al., 2016) to more recent deep learning(Ozsoy, 2016; Vasile et al., 2016; He et al., 2017; Musto et al., 2018), reinforcement learning (Zheng et al., 2018; Shi et al., 2021), language models-based methods (Sun et al., 2019; Fan et al., 2023b). Recently, collaborative filtering-based deep learning techniques gained its popularity back, owing to its simplicity, efficacy and being dependent only on user-item interactions (Mao et al., 2021a; Park et al., 2023). These methods are composed of three major components: interaction module/encoder, loss function and negative sampling strategy (Mao et al., 2021a), The Figure 1 simplifies these steps: Given a representation space, user interacts with an item (e.g., purchases a product), shown with $p$, and from the remaining negative items $n$, four negatives, $n_{1}-n_{4}$, are sampled. Then, the sampled negative items, positive item and user representations are used by the interaction encoding and loss function modules for the computations and making recommendations.
For many years, researchers focused on designing better and more complex interaction modules by utilizing various techniques such as multi-layer perceptrons (Covington et al., 2016; He et al., 2017), auto-encoders (Liang et al., 2018), attention networks (Chen et al., 2020a), transformers (Sun et al., 2019), graph neural networks (He et al., 2020; Ozsoy et al., 2020). Even though these models showed better performance results, they became more and more complex. On the other hand, in industrial recommender systems, highly efficient models are preferred because of their practical applicability.
@@
Recently, SimpleX (Mao et al., 2021a) demonstrated that even a simple interaction encoding model could outperform more complex techniques by utilizing a better-designed loss function. This success has led many researchers in the recommender systems field to focus more on loss functions and negative sampling strategies in the last few years. Several recent studies analyze the impact of loss functions on recommendation performance (Wu et al., 2022; Li et al., 2023; Park et al., 2023). Others (Zhang et al., 2022; Wang et al., 2022; Lee et al., 2023; Sun et al., 2023; He et al., 2023; Wu et al., 2023; Zhang et al., 2024) concentrate on developing new loss functions that address various challenges in recommender systems, such as popularity bias or robustness. These efforts have sparked a growing interest in contrastive learning in recommender systems domain.
@@
The models using contrastive learning aim to learn representations by bringing similar pairs closer while separating dissimilar pairs. The learning process is highly relevant to the designed contrastive loss functions (Wang and Liu, 2021). For an efficient contrastive learning system, three key components have been identified (Weng, 2021): heavy data augmentation, large batch size and hard-negative sampling. While using larger batch sizes and more negative samples introduce more diverse and ‘hard’-negative samples to the system, they require more memory as a side-effect. A common approach is to use a margin parameter to distinguish ‘hard’-negatives from the rest. However, this approach sets only a single threshold, failing to fully utilize all available negative samples efficiently.
@@
In this work, we hypothesise that using not only hardest negative instances but all non-trivial ones allow model to generalize more efficiently, even with smaller number of negative samples or batch sizes. Furthermore, assigning different weights to varying levels of ”hardness” in negative instances—such as giving higher weight to the hardest negatives and relatively lower weights to easier ones—can enhance performance. For this purpose, we propose Multi-Margin Loss (MML), which is a simple yet efficient loss function, that uses multiple margins instead of a single margin and that assigns different weights to different levels of negative instances.
@@
  • Our results show that (i) with a large negative sample size (e.g., 800 negatives), MML performs as well as state-of-the-art contrastive loss functions, and (ii) with smaller sample sizes (e.g., 10 or 100 negatives), MML achieves up to a 20% performance improvement over baseline contrastive loss functions.
The remaining of the paper is structured as follows: Section 2 gives background information on recommender systems, contrastive learning and loss functions. Section 3 presents the proposed Multi-Margin Loss (MML). Section 4 details the experimental setup and presents the evaluation results. Finally, Section 5 discussed the conclusion and future directions.
@@ 2. Related Work
In this section, background information on recommender systems and collaborative filtering, overview of loss functions used in recommender systems and more details specifically on contrastive loss functions are presented.
@@ 2.1. Recommender Systems and Collaborative Filtering-based models
Recommender systems (RS) aims to estimate target users’ preferences and recommend a selected set of items using the historical user-item interaction data. Recommender systems may employ various approaches, such as traditional collaborative filtering, content-based filtering, matrix factorization (Pan et al., 2008; Ye et al., 2010; Li et al., 2015; He et al., 2016) methods or more recent deep learning (Ozsoy, 2016; Vasile et al., 2016; He et al., 2017; Musto et al., 2018), reinforcement learning (Zheng et al., 2018; Shi et al., 2021), language models-based (Sun et al., 2019; Fan et al., 2023b) methods. Recently, collaborative filtering-based deep learning techniques gained its popularity back, owing to its simplicity, efficacy and being dependent only on user-item interactions (Mao et al., 2021a; Park et al., 2023). Collaborative Filtering (CF) models use only user-item interaction history and do not rely on any other contextual information. These techniques are composed of three major components (Mao et al., 2021a; Park et al., 2023); presented in Figure 1:
@@
  • Interaction encoding: Interaction encoder aims to learn embedding of each user and item. It has been the primary focus in the field for many years. To improve it, various techniques such as vector embeddings (Ozsoy, 2016; Vasile et al., 2016), multi-layer perceptrons (Covington et al., 2016; He et al., 2017), auto-encoders (Liang et al., 2018), attention networks (Chen et al., 2020a), transformers (Sun et al., 2019), graph neural networks (He et al., 2020; Ozsoy et al., 2020), reinforcement learning-based approaches (Zheng et al., 2018; Shi et al., 2021) have been employed. While these techniques demonstrated better performance, they became increasingly complex. In industrial recommender systems, however, efficient models are in higher demand due to their practical application (Mao et al., 2021a).
@@
  • Loss function: The commonly used loss functions in recommender systems were Binary Cross-Entropy (BCE) loss (He et al., 2017), Bayesian Personalized Ranking (BPR) loss (Rendle et al., 2009) or mean square error (MSE) loss (Chen et al., 2020b). In the other research domains, such as computer vision, natural language processing, many other loss functions have been developed and applied, which utilise both positive and negative instances and contrastive learning approaches. Inspiring from their success, in the last few years more focus is given to the loss functions in recommender applications.
@@
  • Negative sampling strategy: Negative sampling aims sample subset of negative instances from a larger set (Yang et al., 2024). In recommender systems, there are vast amount of negative or unobserved instances compared to positive ones. In order to improve training efficiency negative sampling techniques are applied. There are various negative sampling strategies, from random negative sampling (RNS) (Pan et al., 2008; Ding et al., 2019), popularity-based negative sampling (PNS) (Barkan and Koenigstein, 2016; Yu et al., 2017; Lian et al., 2020) to hard negative sampling (hard NS) (Zhang et al., 2013; Rendle and Freudenthaler, 2014), and GAN-based negative sampling (GAN-based NS)(Jin et al., 2020).
@@ 2.2. Overview of Loss Functions used in Recommender Systems
  • Point-wise loss function measures the distance between the predicted and the expected labels. Each instance is used independent from the other instances, i.e., either a positive or a negative instance is used at a time.
    Example point-wise loss functions used in recommender systems are Binary Cross Entropy (BCE) loss (He et al., 2017), Mixed-centric (MCL) loss (Gao et al., 2022) and User Interest Boundary (UIB) loss (Zhuo et al., 2022).
  • Pair-wise loss function checks a pair of items at a time and aims to decide the optimal ordering of the pair and compare it to the ground truth. Example pair-wise loss functions used in recommender systems are Bayesian Personalized Ranking (BPR) loss (Rendle et al., 2009), Collaborative Metric Learning (CML) (Hsieh et al., 2017) and Symmetric Metric Learning (SML) (Li et al., 2020).
  • Set-wise loss functions are generalization of the pairwise loss functions (Park et al., 2023). Instead of a pair of positive and negative items, it considers multiple negative items for a single positive item. Example set-wise loss functions used in recommender systems are Cosine Contrastive Loss (CCL) (Mao et al., 2021a), Sampled Softmax (SSM) loss (Wu et al., 2022), Bias-aware Contrastive (BC) loss (Zhang et al., 2022).
Recently, SimpleX (Mao et al., 2021a) demonstrated that well-designed loss function can enable a simple model to outperform more-complex techniques. This success has led to a growing interest in the design of loss functions for recommender systems. Some of the works have analyzed the impact of loss functions on recommendation performance (Wu et al., 2022; Li et al., 2023; Park et al., 2023; Wu et al., 2023). Some others (Zhang et al., 2022; Wang et al., 2022; Lee et al., 2023; Sun et al., 2023; He et al., 2023; Zhang et al., 2024) have focused developing new loss functions to address various challenges, such as popularity bias. These investigations have highlighted contrastive learning as a significant trend that is gaining increasing attention.
@@ 2.3. Contrastive Learning
In Collaborative Filtering, only user-item interactions are used as the input features. Various loss functions, indicated as $f$, are used in the literature for training. For the computations, they use predicted relevance $s(u,x)$, predicted distance $d(u,x)$ and actual label $l(u,x)$ of the user-item interactions. Some of the loss functions use margin $m$ and weight $w$ parameters.
@@
  • Contrastive loss (Chopra et al., 2005): Contrastive loss is one of the earliest loss function defined for contrastive learning. Having pair of user $u$ and item $x$ representation (i.e., $x\in I+$ or $x\in I-$), the goal is to make representation of positive items closer and representation negative items further away from the user representation. The equation for contrastive loss can be written as follows:
    (1) $\begin{split}L_{Contrastive}=\dfrac{1}{2}\sum_{{x}\in{I}}Y_{x}d(u,x)^{2}{\ +}\\ (1-Y_{x})max(0,m-d(u,x))^{2}\end{split}$
    In the equation, $Y_{x}$ indicates function which returns 1 if item is positive (i.e., $x\in I+$) and 0 otherwise. Margin $m$ defines a radius, which makes the model to pay attention to hard-negatives, which are negative instances which has the lowest distance (i.e., most similar) to the user representation. The idea is also presented in Figure 2a. In the figure, margin $m1$ is used for filtering negative samples. As the result, only $n1$, which is in the hard-negatives area, and positive item $p$ are used for the computations.

### v2

@@ Abstract.
Recommender systems guide users through vast amounts of information by suggesting items based on their predicted preferences. Collaborative filtering-based deep learning techniques have regained popularity due to their simplicity, using only user-item interactions. Typically, these systems consist of three main components: an interaction module, a loss function, and a negative sampling strategy. Initially, researchers focused on enhancing performance by developing complex interaction modules with techniques like multi-layer perceptrons, transformers, or graph neural networks. However, there has been a recent shift toward refining loss functions and negative sampling strategies. This shift has increased interest in contrastive learning, which pulls similar pairs closer while pushing dissimilar ones apart. Contrastive learning involves key practices such as heavy data augmentation, large batch sizes, and hard-negative sampling, but these also bring challenges like high memory demands and under-utilization of some negative samples. The proposed Multi-Margin Loss (MML) addresses these challenges by introducing multiple margins and varying weights for negative samples. MML efficiently utilizes not only the hardest negatives but also other non-trivial negatives, offering a simpler yet effective loss function that outperforms more complex methods, especially when resources are limited. Experiments on two well-known datasets showed MML achieved up to a 20% performance improvement compared to a baseline contrastive loss function with fewer negative samples.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Computing methodologies Neural networks
@@ 1. Introduction
Figure 1. Three major components for a deep-learning based CF model: (i) interaction module/encoder, (ii) loss function and (iii) negative sampling strategy. Given a representation space, user interacts with an item (e.g., purchases a product), shown with $p$, and from the remaining negative items $n$, four negatives, $n_{1}-n_{4}$, are sampled. Then, the sampled negative items, positive item and user representations are used by the interaction encoding and loss function modules for the computations.
@@
Numerous web applications, such as YouTube, Amazon, Spotify, use recommendation systems to help users navigate the vast amounts of content (Park et al., 2023). These systems suggest items, such as videos, products, and music tracks, by estimating user preferences (Massa and Avesani, 2007; Tavakolifard and Almeroth, 2012). In the literature, recommender systems employ various approaches, including traditional methods like collaborative filtering, content-based filtering, matrix factorization(Pan et al., 2008; Ye et al., 2010; Li et al., 2015; He et al., 2016) as well as newer techniques such as deep learning(Ozsoy, 2016; Vasile et al., 2016; He et al., 2017; Musto et al., 2018), reinforcement learning (Zheng et al., 2018; Shi et al., 2021), language models-based methods (Sun et al., 2019; Fan et al., 2023b). Recently, collaborative filtering-based deep learning techniques regained popularity due to their simplicity, efficacy and reliance only on user-item interactions (Mao et al., 2021a; Park et al., 2023). These methods are composed of three major components: interaction module/encoder, loss function and negative sampling strategy (Mao et al., 2021a), The Figure 1 simplifies these steps: Given a representation space, user interacts with an item (e.g., purchases a product), shown with $p$, and from the remaining negative items $n$, four negatives, $n_{1}-n_{4}$, are sampled. Then, the sampled negative items, positive item and user representations are used by the interaction encoding and loss function modules for the computations and making recommendations.
@@
For years, researchers focused on improving interaction modules using various techniques such as multi-layer perceptrons (Covington et al., 2016; He et al., 2017), auto-encoders (Liang et al., 2018), attention networks (Chen et al., 2020a), transformers (Sun et al., 2019), graph neural networks (He et al., 2020; Ozsoy et al., 2020). Even though these models showed better performance results, they became more and more complex. On the other hand, in industrial recommender systems, highly efficient models are preferred because of their practical applicability. Recently, SimpleX (Mao et al., 2021a) showed that a simple interaction encoding model can outperform complex techniques with a well-designed loss function. This success has shifted focus in the recommender systems to loss functions and negative sampling strategies. Several recent studies analyze the impact of loss functions on recommendation performance (Wu et al., 2022; Li et al., 2023; Park et al., 2023), while others (Zhang et al., 2022; Wang et al., 2022; Lee et al., 2023; Sun et al., 2023; He et al., 2023; Wu et al., 2023; Zhang et al., 2024) develop new loss functions to address challenges like popularity bias and robustness. These efforts have increased interest in contrastive learning in the recommender systems domain.
@@
The models using contrastive learning aim to learn representations by pulling similar pairs closer while pushing dissimilar pairs apart. The learning process is highly relevant to the designed contrastive loss functions (Wang and Liu, 2021). For an efficient contrastive learning system, three key components have been identified (Weng, 2021): heavy data augmentation, large batch size and hard-negative sampling. While using larger batch sizes and more negative samples introduce more diverse and ‘hard’-negative samples to the system, they require more memory as a side-effect. A common approach is to use a margin parameter to distinguish ‘hard’-negatives from the rest. However, this approach sets only a single threshold, failing to fully utilize all available negative samples efficiently.
@@
In this work, we hypothesise that using not only hardest negative instances but all non-trivial ones allow model to generalize more efficiently, even with smaller number of negative samples or batch sizes. Furthermore, assigning different weights to varying levels of ”hardness” in negative instances—such as giving higher weight to the hardest negatives and relatively lower weights to easier ones—can enhance performance. For this purpose, we propose Multi-Margin Loss (MML), which is a simple yet efficient loss function, that uses multiple margins instead of a single margin and assigns different weights to different levels of negative instances.
@@
  • Our results show that (i) with a large negative sample size (e.g., 800), MML performs as well as state-of-the-art contrastive loss functions, and (ii) with smaller sample sizes (e.g., 10 or 100), MML achieves up to a 20% performance improvement over baseline contrastive loss functions.
The rest of the paper is structured as follows: Section 2 provides background on recommender systems, contrastive learning and loss functions. Section 3 introduces the Multi-Margin Loss (MML). Section 4 covers the experimental setup and evaluation results. Finally, Section 5 discusses the conclusion and future directions.
@@ 2. Related Work
This section provides background on recommender systems and collaborative filtering, an overview of loss functions used in these systems, and more details on contrastive loss functions.
[…]
@@ 2.1. Recommender Systems and Collaborative Filtering-based models
Recommender systems (RS) aim to estimate users’ preferences and recommend items based on historical user-item interaction data. RS can use various approaches, such as traditional collaborative filtering, content-based filtering, matrix factorization (Pan et al., 2008; Ye et al., 2010; Li et al., 2015; He et al., 2016), deep learning (Ozsoy, 2016; Vasile et al., 2016; He et al., 2017; Musto et al., 2018), reinforcement learning (Zheng et al., 2018; Shi et al., 2021), and language model-based methods (Sun et al., 2019; Fan et al., 2023b). Recently, collaborative filtering-based deep learning techniques have regained popularity due to their simplicity, efficacy, and reliance only on user-item interactions (Mao et al., 2021a; Park et al., 2023). Collaborative Filtering (CF) models use only interaction history without any other contextual information and consist of three major components (Mao et al., 2021a; Park et al., 2023), as shown in Figure 1.
@@
  • Interaction encoding: Interaction encoder aims to learn embedding of each user and item. It has been primary focus in the field for many years. To improve it, various techniques such as vector embeddings (Ozsoy, 2016; Vasile et al., 2016), multi-layer perceptrons (Covington et al., 2016; He et al., 2017), auto-encoders (Liang et al., 2018), attention networks (Chen et al., 2020a), transformers (Sun et al., 2019), graph neural networks (He et al., 2020; Ozsoy et al., 2020), reinforcement learning techniques (Zheng et al., 2018; Shi et al., 2021) have been employed. Despite better performance, these methods became increasingly complex. In industry, however, efficient models are in higher demand due to their practical application (Mao et al., 2021a).
@@
  • Loss function: The commonly used loss functions in RS were Binary Cross-Entropy (BCE) loss (He et al., 2017), Bayesian Personalized Ranking (BPR) loss (Rendle et al., 2009) or mean square error (MSE) loss (Chen et al., 2020b). In the other research domains, such as computer vision, natural language processing, many other loss functions have been developed and applied, which utilise both positive and negative instances and contrastive learning approaches. Inspiring from their success, in the last few years more focus is given to the loss functions in recommender applications.
@@
  • Negative sampling strategy: Negative sampling aims sample subset of negative instances from a larger set (Yang et al., 2024). In RS, there are vast amount of negative or unobserved instances compared to positive ones. In order to improve training efficiency negative sampling techniques are applied. There are various negative sampling strategies, from random negative sampling (RNS) (Pan et al., 2008; Ding et al., 2019), popularity-based negative sampling (PNS) (Barkan and Koenigstein, 2016; Yu et al., 2017; Lian et al., 2020) to hard negative sampling (hard NS) (Zhang et al., 2013; Rendle and Freudenthaler, 2014), and GAN-based negative sampling (GAN-based NS)(Jin et al., 2020).
@@ 2.2. Overview of Loss Functions used in Recommender Systems
  • Point-wise loss function measures the distance between the predicted and the expected labels. Each instance is used independent from the other instances, i.e., either a positive or a negative instance is used at a time. Example point-wise loss functions used in RS are Binary Cross Entropy (BCE) (He et al., 2017), Mixed-centric (MCL) (Gao et al., 2022) and User Interest Boundary (UIB) (Zhuo et al., 2022) losses.
  • Pair-wise loss function checks a pair of items at a time and aims to decide the optimal ordering of the pair and compare it to the ground truth. Example pair-wise loss functions used in RS are Bayesian Personalized Ranking (BPR) loss (Rendle et al., 2009), Collaborative Metric Learning (CML) (Hsieh et al., 2017) and Symmetric Metric Learning (SML) (Li et al., 2020).
  • Set-wise loss functions are generalization of the pairwise loss functions (Park et al., 2023). Instead of a pair of positive and negative items, it considers multiple negative items for a single positive item. Example set-wise loss functions used in RS are Cosine Contrastive Loss (CCL) (Mao et al., 2021a), Sampled Softmax (SSM) (Wu et al., 2022), Bias-aware Contrastive (BC) (Zhang et al., 2022) losses.
Another direction in recommender systems is the use of automated machine learning (AutoML) techniques to explore loss functions (Chen et al., 2024). SFL (Liu and Lai, 2020) and AutoLoss (Zhao et al., 2021) aim to combine loss functions automatically, while AutoLossGen (Li et al., 2022) generates new loss functions from basic mathematical operators. We believe that creating well-designed loss functions and applying AutoML techniques for their generation and exploration will complement each other, enhancing the overall performance of recommender applications.
Recently, SimpleX (Mao et al., 2021a) demonstrated that a well-designed loss function can enable a simple model to outperform more complex techniques, sparking interest in designing loss functions for recommender systems. Some studies have analyzed the impact of loss functions on recommendation performance (Wu et al., 2022; Li et al., 2023; Park et al., 2023; Wu et al., 2023), while others (Zhang et al., 2022; Wang et al., 2022; Lee et al., 2023; Sun et al., 2023; He et al., 2023; Zhang et al., 2024) have developed new loss functions to address challenges like popularity bias. These investigations have highlighted contrastive learning as an increasingly popular trend.
@@ 2.3. Contrastive Learning
In Collaborative Filtering, only user-item interactions are used as the input features. Various loss functions, indicated as $f$, are used in the literature for training. For the computations, they use predicted relevance $s(u,i)$, predicted distance $d(u,i)$ and actual label $l(u,i)$ of the user-item interactions. Some of the loss functions use margin $m$ and weight $w$ parameters.
@@
  • Contrastive loss (Chopra et al., 2005): Contrastive loss is one of the earliest loss function defined for contrastive learning. Having pair of user $u$ and item $i$ representation (i.e., $i\in I+$ or $i\in I-$), the goal is to make representation of positive items closer and representation negative items further away from the user representation. The equation for contrastive loss can be written as follows:
    (1) $\begin{split}L_{Contrastive}=\dfrac{1}{2}\sum_{{x}\in{I}}Y_{i}d(u,i)^{2}{\ +}\\ (1-Y_{i})max(0,m-d(u,i))^{2}\end{split}$
    In the equation, $Y_{i}$ indicates function which returns 1 if item is positive (i.e., $i\in I+$) and 0 otherwise. Margin $m$ defines a radius, which makes the model to pay attention to hard-negatives, which are negative instances which has the lowest distance (i.e., most similar) to the user representation. The idea is also presented in Figure 2a. In the figure, margin $m1$ is used for filtering negative samples. As the result, only $n1$, which is in the hard-negatives area, and positive item $p$ are used for the computations.
@@
    AdvInfoNCE(Zhang et al., 2024) further improves InfoNCE loss by adaptively exploring and assigning a difficulty level to each negative instance in an adversarial manner.